Journalist Report – February 12th

Crew 205 Crew Journalist Report 12-FEB-2019

Sol 2

Author: Maria Grulich

Title: EVAs EVAs…

Second day on Mars we woke up with the smell of scrambled eggs and spam and self-made brownies that we prepared last night.

As a lesson learned we started to prepare earlier today for the morning EVA to collect rock samples.

The EVA 1 crew: Nathan, Daniel, Natalia and Maria left the hab at 0930AM. First, EVA1 crew took some pictures with signs for outreach before heading to the White Moon area using the rovers Spirit and Opportunity.

The road was windy, and every corner revealed a new area of Mars. Opportunity rover was doing great whereas the Spirit rover concerned EVA1 crew more, as the State of Charge seems to be dropping fast. Therefore, EVA1 had to stop earlier than planned to start walking by foot and leave the rovers behind.

Finding the White Moon seems to be more challenging than earlier EVAs as the road was not properly marked and the helmets of the space suits kept fogging and limiting visibility. EVA1 lost communication with the hab because they were out of reach of the communication system, which added more challenges.

During this walk the blazing Sun revealed shimmers and flashes of light from the crews´ visors that were actually shiny rocks (believed to be quartz containing crystals) underneath their feet. They stopped to retrieve these jewels in the desert as their first sample of the day.

EVA1 crew kept on walking, stopping from time to time to figure out where they actually were until they arrived at their desired location: White Moon. The white soil of the hill they stopped on let us guess why they called it the Moon. Small stains of red underneath the chalk-white hill intrigued the crew, and a short hit with the hammer revealed a bright red soil underneath. Described as a geological “red velvet cake”, it was great for their samples.

The curiosity kept running through EVA1 crew veins and let them keep moving anuntil they found crevice. Everyone of that EVA got so excited as this indicated that there had been water running through this land a long time ago. Two crew members climbed down into the crevice to take samples of the different sediments.

While EVA1 were busy in the White Moon, the rest of the crew were busy working in the hab. The crew astronomer was busy working on photometry and setting up new observation for tonight. Work in the GreenHab was also ongoing by watering the plants and making sure everything is in place.

Pumped up with excitement of their discoveries, the crew went back towards the hab where the support crew welcomed them with a hot chicken soup.

Just one hour later after EVA1 crew arrived EVA2 crew was about to get ready. During the preparation it was clear that not all space suits are working properly, and we are unclear on how long they have to be charged before being used again.

As there were not enough spare suits that not had been used before, or where the fans are running properly, the commander came to the decision that the EVA2 has to be scrubbed. It was a tough call but Safety first.

After that first difficult decision the team meeting was all about how to improve the EVA preparation to achieve two EVAs a day that would accommodate all our science objectives. During the analysis of the first rock samples in the science dome it was observed that the state of charge of the solar panels dropped significantly and that was a wakeup call for everyone:

Mars is not easy it is hard.

We have to keep more focus on the maintenance of the habitat while being focused on our personal goals. Still the crew found time for their first yoga sessions and preparing for the Martian Spanish night with tapas and Spanish jamón.

Tomorrow we will be planning for two EVAs again and making sure to keep checking and preparing accordingly.

Ad Astra!

Journalist Report – November 25th

Crew 205 Crew Journalist Report 12-FEB-2019

Sol 2

Author: Maria Grulich

Title: EVAs EVAs…

Second day on Mars we woke up with the smell of scrambled eggs and spam and self-made brownies that we prepared last night.

As a lesson learned we started to prepare earlier today for the morning EVA to collect rock samples.

The EVA 1 crew: Nathan, Daniel, Natalia and Maria left the hab at 0930AM. First, EVA1 crew took some pictures with signs for outreach before heading to the White Moon area using the rovers Spirit and Opportunity.

The road was windy, and every corner revealed a new area of Mars. Opportunity rover was doing great whereas the Spirit rover concerned EVA1 crew more, as the State of Charge seems to be dropping fast. Therefore, EVA1 had to stop earlier than planned to start walking by foot and leave the rovers behind.

Finding the White Moon seems to be more challenging than earlier EVAs as the road was not properly marked and the helmets of the space suits kept fogging and limiting visibility. EVA1 lost communication with the hab because they were out of reach of the communication system, which added more challenges.

During this walk the blazing Sun revealed shimmers and flashes of light from the crews´ visors that were actually shiny rocks (believed to be quartz containing crystals) underneath their feet. They stopped to retrieve these jewels in the desert as their first sample of the day.

EVA1 crew kept on walking, stopping from time to time to figure out where they actually were until they arrived at their desired location: White Moon. The white soil of the hill they stopped on let us guess why they called it the Moon. Small stains of red underneath the chalk-white hill intrigued the crew, and a short hit with the hammer revealed a bright red soil underneath. Described as a geological “red velvet cake”, it was great for their samples.

The curiosity kept running through EVA1 crew veins and let them keep moving anuntil they found crevice. Everyone of that EVA got so excited as this indicated that there had been water running through this land a long time ago. Two crew members climbed down into the crevice to take samples of the different sediments.

While EVA1 were busy in the White Moon, the rest of the crew were busy working in the hab. The crew astronomer was busy working on photometry and setting up new observation for tonight. Work in the GreenHab was also ongoing by watering the plants and making sure everything is in place.

Pumped up with excitement of their discoveries, the crew went back towards the hab where the support crew welcomed them with a hot chicken soup.

Just one hour later after EVA1 crew arrived EVA2 crew was about to get ready. During the preparation it was clear that not all space suits are working properly, and we are unclear on how long they have to be charged before being used again.

As there were not enough spare suits that not had been used before, or where the fans are running properly, the commander came to the decision that the EVA2 has to be scrubbed. It was a tough call but Safety first.

After that first difficult decision the team meeting was all about how to improve the EVA preparation to achieve two EVAs a day that would accommodate all our science objectives. During the analysis of the first rock samples in the science dome it was observed that the state of charge of the solar panels dropped significantly and that was a wakeup call for everyone:

Mars is not easy it is hard.

We have to keep more focus on the maintenance of the habitat while being focused on our personal goals. Still the crew found time for their first yoga sessions and preparing for the Martian Spanish night with tapas and Spanish jamón.

Tomorrow we will be planning for two EVAs again and making sure to keep checking and preparing accordingly.

Ad Astra!

Sol Summary – February 12th

Crew 205 Sol Summary Report 12- FEB – 2019

Sol: 2

Summary Title: Adapting to Mars

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: We started the day with what is becoming our daily routine: waking up at 7 am, having coffee and breakfast together and by 8.30 am start our work. In the morning, the first EVA team set off on rovers Spirit and Opportunity Northwards on Cow Dung Road towards the “Moon” areas. They passed Galileo road (losing contact with MDRS as expected shortly after). Before reaching Cowboy Corner, Spirit reported having 75% SOC, Spirit’s battery reduced rapidly and the rovers were stopped between Tank Wash and White Moon with Spirit on 63% and Opportunity at 75%. The EVA was modified to be completed on foot by only visiting, exploring and sampling areas within White Moon and the easternmost fringes of Beige Moon. The turnings of the road that enter to the Moon were recorded on GPS for future use in EVAs. Some excellent samples of quartz and sulphur containing deposits were collected and will be analysed in the coming days before being described fully in the upcoming Science Report. Meanwhile, in the Hab, the rest of the team continue working in the GreenHab daily duties, EVA planning and astronomy projects. Lunch was ready for the returning EVA team and the crew ate together in good spirits.

Equipment was prepared for use in the second EVA of the day, including a drone with a vertical camera and the ability to connect to a laptop for images to be reviewed in the field. As planned, the EVA team started getting prepared for its mission to North Ridge. However, while preparing, it was found that some of the suits were not fully functional and others, were still charging from the EVA in the morning. The team initially thought only of the suits was to be fully charged and operational. For this reason, putting the safety of the crew as the top priority, the commander took the decision to cancel the EVA. After the fact, it was found that a second spacesuit was also fully charged. All crew members thus, stayed in the Hab working “from home” in their different duties and projects. Tonight, following our daily crew meeting, we plan to enjoy a “Martian-Spanish” tapas dinner.

Look Ahead Plan: Tomorrow we plan to continue with the initial plan of today. A first EVA team will head in the morning will head to Maxwell Montes. A second team will head the Old Repeater Point and North Ridge to test the drone. GreenHab, astronomy and outreach activities will continue as usual.

Anomalies in work: EVA#04 (in the afternoon) was cancelled (please refer to EVA report and description above for further information)

Weather: Clear skies all day long

Crew Physical Status: Nominal

EVA:

EVA#03: 4254500/516500 (Gray Moon, not reached), 4254350/516500 (Beige Moon), 4254550/517550 (White Moon)

EVA#04: No conducted (please refer to the EVA report and description above for more information)

Reports to be filed: Sol Summary, EVA request, Operations Report, Astronomy Report, Greenhab Report, Journalist Report

Support Requested: Some spacesuits not working properly and we are missing a charger for one of them (please refer to operations report for further information).

Operations Report – February 12th

Crew 205 Operations Report 12 – FEB – 2019

SOL: 2

Name of person filing report: Veronica Trivino

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on this evening by 1708 (12thFeb), Stopped by this morning 0730 (12thFeb)

Solar— SOC % – Turned off Gen (12th Feb) 33%, Turned on Gen (12th Feb) 69%

Diesel Reading – 100%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 77%, ~ 422 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 276.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01411709 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49.6 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used

Hours: 73.7

Beginning charge: 100%

Ending charge: 40%

Currently charging: Yes

Notes on rovers: Spirit was accidentally left on after returned from EVA no.1. Engineer when checking readings in the evening turned it off. It was charged to 100% when checked by engineer.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? Not used

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Several suits with fan broken. Other suits had not finished charging by the time the second EVA preparation had begun, even though they were set to charge as soon as the crew from the first EVA came back. One of the suit chargers is not functional. More information was sent to Atila/Mission support in an email.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Greenhab Report – February 12th

Crew 205 Green Hab Report – 12-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (4 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.7°C

High: 27.9°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 276.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%) [AT TIME OF WATERING]

1) 09.00 AM 19.5° C / 54 %
2) 12.00 PM 27.6° C / 28 %
3) 15:30 PM 19.8° C / 63 %
4) 18:15 PM 18.7° C / 45 %

Changes to crops: None

Narrative: Today we continued sample collection on EVA’s for the regolith samples for the nutrient supplementation experiments (see mission summary). We also began characterizing regolith samples and took photographs underneath the dissection microscope. A more thorough analysis of the samples will be available in tomorrow’s science report. The A. Thaliana seeds were washed and placed on germination plates last night and we expect germination to be in initial stages tomorrow.

Additionally, we took photographs of the dissected cucumber samples under the optical microscope and found evidence of degraded cell structure, likely from the fungi growing on the dead plants. More information and photos will be available in tomorrow’s science report as well.

We removed 1 dead sunflower from its grow bed to prevent the spread of mold. Similarly, we removed 1 dead Red Leaf Lettuce from its pot. Additionally, we groomed dead biomass (fallen leaves/stems) from several of the grow beds. We also harvested small amounts of oregano, parsley and lemon basil for lunch.

Harvest: 1/3 g Oregano

1/3 g Parsley

1/3 g Lemon basil

6.3 g of dead material (including dead cucumbers)

Support/supplies needed: None.

Astronomy Report – February 12th

Crew 205 Astronomy Report 12 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-14

Objects Viewed: ASASSN-V J081823.00-111138.9. I called it MDRS Target 1 in Skynet (RA: 08 18 23.00 , DEC: -11 11 38.9). The observations are planned to be 120 s exposure time with the B filter and 60 s with the V filter. 9 exposures were requested for a period of about 3 hours. A photometry value was calculated for the target SY MON (that was observed in 11 Feb 2019), and the photometry value was found to be 14.147. This value will be submitted to the AAVSO.

Problems Encountered: Nothing to mention

Journalist Report – February 11th

Crew 205 Crew Journalist Report 11-FEB-2019 Sol 1

Author: Maria Grulich

Title: One small step…

Today we woke up on Mars! The weather looks great, the Red Planet is really showing us its best side. We are just not sure what the white crystal soil is outside…

The morning was preparing for our first crew EVA: Daniel (HSO), Dave (XO), Hannah (COMMS) and Nathan (GHO) are ready to step on Mars to explore the wonders of the Red Planet. Due to the unexpected white crystal soil, crew 205 was instructed to a different location closer located to the hab, near the rock formation known as the “Kissing Camel”. Maria (Media) was in the role of Hab Control as EVA support.

The EVA started with getting into the rovers which seems to be challenging to get used to, due to the size of the space suit (described as “a mid-western barn” by our XO). Once at the location, the new habitants of Mars collected their first samples. These samples will be used later for growing plants in the base’s GreenHab. Communication between Hab and EVA1 crew was difficult due to issues with microphones and distortion.

The second EVA at 1500 with Natalia (Commander), Ghanim (Astro), Veronica (Engineer) and Maria (Media) was supported by Daniel and Hannah as Hab Control. The sun was shining and walking on the red planet was a lot harder than expected. The EVA2 crew members collected another sample.

Earlier than expected the crew experienced rising temperatures in their suits and increasing fog affecting the view. The EVA2 crew therefore decided to return early to avoid overheating. Walking on Mars seems to be easier than it actually is.

The support crew was ready to welcome them with water to cool everyone down.

The lesson for the following days is to reduce the number of layers EVA crews wear, as due to the additional weight of the space suits the body temperature is raised naturally and less warm clothing is needed.

The day was closed with our first dinner on Mars serving mashed potatoes, spam and the first harvested carrots of the Green Hab!

We are very excited for the days to come and hope that we keep the spirit up!

Ad Astra!

Operations Report – February 11th

rew 205 Operations Report 11 – FEB – 2019

SOL: 1

Name of person filing report: Veronica Trivino

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1730 (11th Feb), Stopped by this morning 0730 (11th Feb)

Solar— SOC % – Turned off Gen (11th Feb) 21%, Turned on Gen (11th Feb) 70%

Diesel Reading – 25%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 88%, ~ 485 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 288.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01411118 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49.2 hours

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used

Hours: 73.2

Beginning charge: 100%

Ending charge: 85%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, Off campus

Crew Car used and why, where? Not used

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Suit 10 – Fan broken. Email sent to to Scott for repair aid.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: We were operating the autoclave by the handbook that was provided and as the pressure began to rise to the optimal pressure, it began to drop suddenly. We found that there was a leak out of the threads of the pressure valve that attaches the whole valve assembly to the lid, probably because it was not tightened properly. In order to remedy that, the XO got an adjustable wrench and made a tightening quarter turn on the valve, which caused the leak to stop and pressure, began to rise again. However, a minute later, the upper pressure release segment popped off. Upon inspection, we found that the valve sheared where pressure normally releases. It appears that the valve was already weakened as it sheared cleanly. No threads were damaged, so if a replacement valve assembly can be obtained, the issue should be repairable. The minimum sterilization pressure (17 psi) was obtained only momentarily and no higher pressures were experienced.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: EVA 2 encountered 2 cases of overheating crew, one light, one more significant. All treated and crew were healthy again within 25 mins. See more in HSO report.

Questions, concerns and requests to Mission Support: None

Sol Summary – February 11th

Sol: 1

Summary Title: First day

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: Today was the first day of sim. We woke up looking at the snowy Martian landscape from our small windows in the Hab. Slowly, we are adapting to the life here. We are getting familiar with our new home and our new crew family. We began the day by removing the snow from the ground in the tunnels and enjoying breakfast all together. Today we had the opportunity to experience our first EVAs. The snow however, introduced some changes to our initial plans. The first EVA team scheduled in the morning, originally planning to reach the White Moon area, was redirected to Kissing Camel Ridge. The team conducted some sample collection and explored the area. The second EVA team in the afternoon headed to North Ridge. Due to the hot weather, the team came back earlier than expected. We learned the challenges of moving around with the spacesuit under the “Martian” sun. Meanwhile, we started our project on food production in regolith; although it was challenged with some issues with the autoclave. We harvested some of the carrots from the GreenHab. We also performed our first astronomical observation (we expect to download the data later today).

In the evening, we did inventory of the food that we received today (shipped form “Earth”), arranged the kitchen and held our daily crew meeting. Tonight, a nice dinner including those Martian harvested carrots is waiting for us!

Looking Ahead Plan: Tomorrow we plan to also conduct two EVAs following the original plans scheduled for today: one team will head to The Moon area; the second team will head to North Ridge. We will test the drones for our EVA navigation projects, continue regolith sample collections, and continue working on the other research projects and duties (including photometry astronomical analysis, GreenHab and outreach activities).

Anomalies in work: None

Weather: Light snow on the ground in the morning. Clear skies all day long.

Crew Physical Status: Nominal

EVA:

EVA#01: 4249300,518100

EVA#02: 4250500, 518400

Reports to be filed: Sol Summary, EVA request, EVA Reports, Operations Report, Greenhab Report, Journalist Report, HSO report, Mission plan, Astronomy Report.